# Future proofing network encryption technology (and securing critical infrastructure data) with continuous-variable quantum key distribution

### v1

@@
Abstract
@@
We demonstrate the establishment of quantum-secure data transfer links at two locations in Denmark: on the campus of Technical University of Denmark (DTU) in Lyngby and between two power grid nodes owned and operated by Energinet in Odense. Four different channels were investigated, one being a purely underground fiber and the other three being combinations of underground fibers and optical ground wires (OPGWs). Coherent ‘quantum’ states at 1550 nm, prepared and measured using a semi-autonomous continuous-variable quantum key distribution (CVQKD) prototype, were multiplexed in wavelength with ‘classical’ 100G encrypted data traffic from a pair of commercial layer-2 network encryption devices operating at around 1300 nm. At DTU, we estimate average secret key rates in the asymptotic limit of $1.12$ Mbps (channel loss of 5.5 dB at 1550 nm) while at Energinet, the figures corresponding to the three channels (with losses of 4.1, 6.7, and 8.9 dB) are $2.05$, $0.90$, and $0.23$ Mbps, respectively. The demonstration showcases that QKD can serve as an additional layer to protect sensitive network traffic propagating on insecure channels.
I Introduction
@@
The explosion in the data traffic volumes in the last few years, driven by rapid growth of data centers and cloud service providers, poses significant security challenges. This is due to the nature of the data involved, which ranges from critical national infrastructure and user identities to financial secrets such as intellectual property, all of which are at risk. In particular, the energy infrastructure has been increasingly targeted, as evidenced by incidents such as the breach at a nuclear power plant in India [1], the Colonial Pipeline ransomware case in USA [2], and most recently, the coordinated hacking of 22 different energy companies in Denmark [3]. These events highlight a worrisome trend of increasing frequency, sophistication, and involvement of nation-states in cyberattacks.
@@
While implementing good practices, such as timely application of patches and measures to prevent exploitation of vulnerabilities can never be underscored more, cryptographic solutions such as encryption add a protective layer on the sensitive data to foil adversaries. In the industrial context, a typical approach involves the use hardware encryption devices, also known as network encryptors or Ethernet encryptors [4]. The devices, employing symmetric cryptographic algorithms such as advanced encryption standard (AES), operate in pairs, utilizing the same sequence of pre-configured keys for encryption and decryption of data.
@@
A natural method to distribute random symmetric keys across an insecure communication channel is provided by quantum key distribution (QKD) [5, 6, 7]. In QKD, two users, Alice and Bob, can share quantum correlations through the exchange of quantum-optical signals, while ensuring that an eavesdropper, Eve, cannot intercept these correlations without getting disclosed. This is because any attempt by Eve to eavesdrop introduces noise and/or loss that Alice and Bob can quantitatively estimate, assuming pre-calibrated devices.
@@
In this work, we showcase an example of quantum-classical integration for securing data in transit using QKD and network encryption. A semi-autonomous, telecom-rack mountable continuous-variable (CV) QKD system [13, 14, 15], operating at 1550 nm, produced a secret keystream for a pair of layer 2 network encryptors [16], communicating simultaneously over the same channel using 100G optical transceivers at 1300 $\pm$ 10 nm.
Figure 1: Practical network encryption with continuous-variable quantum key distribution. (Top) Map of Denmark in the center, with the geographical locations for the experiments indicated using the cyan and yellow pins. On the left is a satellite view showing the power substation locations in Odense, and on the right is a zoomed-in map of DTU. (Bottom) Optical connections are depicted using dashed-green, dotted-blue or solid-orange lines depending on the wavelength(s) of light traversing through them, with a single/pair of line(s) indicating simplex/duplex. A fixed attenuator and a pair of patchcords provide a path for signals at $\sim$ 1300 nm to traverse from RX to TX. The CVQKD-TX and RX modules comprise a solid plate with the optical/optoelectronic components on one side (as shown) and the necessary control electronics on the other. The layer 2 encryptors utilize the AES encryption algorithm with 256 bit long keys. The main text contains a detailed description of all relevant devices as well as the system operation. DPI: digital processing interface.
Experiments were conducted at two locations in Denmark using four different combinations of underground and aerial fibers acting as the channel for the propagation of the C-band quantum states and the O-band optical signals. Off-the-shelf coarse wavelength division multiplexers (CWDMs) were used for their union and separation. The experimental scheme, illustrated in Figure 1, involved initial tests on a loopback single-mode fiber (SMF) under the Lyngby campus of Technical University of Denmark (DTU) to optimize the system operation, followed by temporary installation and testing at the premises of Energinet in Odense. Here, we explored three different loopback topologies using optical ground wires (OPGWs) running between power substations located at Fraugde and South East Odense (Odense SØ), as indicated in the top-left map in Fig. 1. More specifically, two of the links (L2 and L4) were a 2$\times$ and 4$\times$ back-and-forth loop of these aerial fibers, while L3 was a combination of an OPGW and an underground SMF.
Table 1 summarizes the link lengths, losses (at 1550 nm), and the final results—secret key fractions (SKFs) obtained in the asymptotic limit—for these links. Assuming real-time signal and data processing at the operating rate ($B=20\,$MBaud) of our CVQKD system, multiplying the SKF by $20\times 10^{6}$ yields the achievable secret key rate (SKR). With this assumption, the number of AES 256-bit secret key pairs for any of the investigated links far exceeds the key-pair refresh rate requirement of 30 per hour for continuous and real-time operation of the network encryptors.
The CVQKD transmitter (CVQKD-TX) and receiver (CVQKD-RX) are standard 3U 19-inch modules suitable for telecom rack installation.
Link label Location Length, km Physical loss, dB SKF, bits/symbol
L1
Lyngby
$0.55$
$5.5$
$0.06\pm 0.02$
L2
Odense
$13.2$
$4.1$
$0.11\pm 0.02$
L3
Odense
$15.3$
$6.7$
$0.05\pm 0.02$
L4
Odense
$25.9$
$8.9$
$0.01\pm 0.01$
@@
Table 1: Information about the links investigated and the secret key fraction (SKF) obtained thereon. All 4 links were loopbacks in the sense that both the input and output fiber were physically accessible at the same place. The loss and length at Lyngby were estimated using simple optical power and pulse delay measurements. At Odense, these quantities were determined using an optical time domain reflectometer.
This CVQKD system, coined qTReX [17], uses randomness from a state-of-the-art quantum random number generator (QRNG) [18] and a host of digital signal processing (DSP) methods [19, 13, 15] to enable the sharing of quantum correlations between Alice and Bob. Classical information processing steps, applied offline on sequences of data frames from the CVQKD-TX and -RX, utilize these correlations to distill secret keys for the encryptors. The balanced coherent detection employed in CVQKD-RX, together with the large wavelength separation (1550 nm vs 1300 nm) of the quantum and classical signals, ensures that despite being more than 7 orders of magnitude higher in launch power, the classical signal does not have any observable influence on the quantum signal noise characteristics observed at CVQKD-RX.
Currently, the entire setup can operate on the physical layer with minimal user intervention. Once the classical data processing stack is also brought online, the entire operation from quantum states’ production to live data encryption would become fully autonomous. With a few hardware modifications, the qTReX system can be adopted in data-center-like environments [20]. Our work thus serves as a proof-of-concept demonstration of using QKD to ‘future proof’ network encryption technology. We note that QKD has been previously used in a deployed electric utility fiber network for secure authentication of smart grid communications [21].
@@
The paper is organized as follows: In the next section, we provide details of the complete quantum-classical setup that enables future proofing network encryption with CVQKD. This is followed by a description of the conducted experiments and the obtained results. In particular, we describe a novel aspect of the implementation, namely the live optimization of the state of polarization (SOP) of the 1550 nm signal, which is crucial for obtaining a good performance during the quantum phase of the QKD protocol. Finally, we discuss some of the results and provide an outlook before concluding the paper.
II Scheme: Channels & devices
@@
In this section, we present the building blocks of the scheme illustrated in the center of Fig. 1. More specifically, we elaborate on the different communication channels, the connected devices, and their roles in the different stages of the protocol. For an introduction to the principles of CVQKD as well as typical CVQKD schemes, we refer the reader to reviews in Refs. [6, 22, 23, 24].
II.1 Quantum channel
@@
However, the quantum-optical signal coded at the sideband frequencies is relatively dim. To address the tasks of carrier and clock recovery—necessary for phase-reference alignment and synchronization of the RX and TX—the CVQKD system utilized some ancillary signals with higher signal to noise ratio (SNR). These signals are described later in this section. In contrast to a typical lab-based setup that often uses electronic signals for triggering and reference-clock-sharing between the RX and TX, here, the sole signal path was a one-way optical connection. This is the quantum channel: the thick orange line and the dotted blue lines joined together using the CWDMs in Fig. 1, with optical waveforms labelled $TX_{\text{out}}$ and $RX_{\text{in}}$ (bottom of the figure) further visualizing the entry and exit points of this channel.
II.2 Classical channel
While the CVQKD devices were connected solely through the quantum channel, their respective digital processing interfaces (DPIs) were connected ‘classically’. Specifically, a pair of workstations on the local network formed a classical channel (not shown in Fig. 1) for the exchange of data between TX and RX during the classical phase of the QKD protocol. Ideally, this classical channel should be authenticated [6], typically achieved using a part of the secret key from an earlier round of the protocol. However, to simplify the overall scheme, we did not implement such authentication.
II.3 Encryption channel
For high-speed network encryption (in the 1-100 Gbps range) over dedicated point-to-point links serving data centers, cloud service providers, and large enterprises, fiber-optic small form factor pluggables (SFPs) are the standard solution for distances spanning several kilometers [26].
We used off-the-shelf 100G QSFP28 4WDM-40 hot-pluggable modules [27] for transmitting and receiving the encrypted data. As depicted by the dashed-green line pairs in Fig. 1, these transceiver modules operate in duplex mode. To ensure bidirectional traffic at 1300 $\pm$ 10 nm without saturating the SFP receiver at the encryptor on the TX side, we inserted a fixed attenuator (loss $\sim$ 5 dB at 1300 nm) between the CWDMs. The encryption paths thus constituted the quantum channel and fiber patch cords with the fixed attenuator (TX to RX: solid-orange line, RX to TX: dashed-green line).
II.4 CVQKD-TX
@@
The flowchart in Fig. 2(a) illustrates the DSP recipe for creating the digital input to the DAC, while the spectrum of the complex waveform, whose real (imaginary) part is related to RF1 (RF2), is depicted in Fig. 2(b).
Figure 2: Digital signal processing (DSP) in qTReX. (a) Flowchart showing the various DSP steps for encoding of the quantum states in a $B=20\,$MHz wide sideband together with several time- and frequency-multiplexed ancillary signals. (b) Spectrum of the complex signal, the real and imaginary parts of which are played out using the two channels of the DAC. (c) Spectra of the ADC captured signals under various configurations of the optical switches in CVQKD-RX, detailed further in the text. Vac: quantum vacuum state $|0\rangle$, LO: local oscillator, Q/PRNG: quantum/pseudo random number generator. CAZAC: constant amplitude zero auto correlation.
Uniformly distributed random numbers from the QRNG were transformed into discrete Gaussian-distributed numbers with 6 bits of resolution and a range of 7 standard deviations [18]. To this quantum data, a constant amplitude zero auto correlation sequence was multiplexed in time, while a quadrature phase shift keying (QPSK) data sequence containing a known header and frame identification (ID) information was multiplexed in frequency [15]. The subsequent step of pulse shaping was performed by upsampling to the DAC sampling rate of 1 GSamples/s and applying root raise cosine filtering [13, 14]. After single sideband modulation of the two 20 MHz wide sidebands at 80 and 160 MHz, two pilot tones at 25 MHz and 120 MHz were also multiplexed in frequency as depicted in Fig. 2(b).
Referring to Fig. 1 (bottom), the optical power at the output of the IQ modulator was monitored using a 1:99 fiber splitter and power meter. We could independently set the 1) modulation strength associated with the quantum states, 2) power of the pilot tones, and 3) power in the QPSK data band to desired levels through a combined digital control of signal amplitudes during the DSP procedure, tuning of the voltage output levels of the DAC, and the optical attenuation using an electronically controlled variable attenuator after the IQ modulator.
All fiber-optic components featured polarization maintaining fiber (PMF) pigtails, except the SMF-connectorized wavelength filter and isolator (added to prevent Trojan-horse attacks [29, 30]). The signal $\text{TX}_{\text{out}}$ was fed by a simplex cable to the input of CWDM-TX.
II.5 CVQKD-RX
@@
Due to arbitrary polarization fluctuations in the channel, the incoming signal $\text{RX}_{\text{in}}$ required polarization control; section III.1 describes the details of this procedure. The signal polarization is in a desired state when it aligns with the polarization of the LLO, where after, the ADC captured the modulated signal.
@@
II.6 Digital Processing Interface
Two workstations formed the DPIs for operating the entire QKD protocol. PCI Express based DAC and ADC cards, housed in these workstations, were used for driving the modulator in CVQKD-TX and acquiring the detector output in CVQKD-RX, respectively, as explained above. The DSP cycles running on each of them allowed the quantum symbols encoded at CVQKD-TX to be eventually decoded at CVQKD-RX, i.e., fulfil the state preparation and measurement tasks. Figure 3 in Ref. [15] shows a flowchart of the DSP running on CVQKD-RX, complementary to the one Fig. 2(a).
Thereafter, graphic processing unit (GPU) equipped workstations communicated over the classical channel to perform the remaining data processing stages of the protocol, namely information reconciliation (IR), parameter estimation (PE), and privacy amplification (PA). Usage of GPU speeds up the execution of both IR and PA.
II.7 Network encryptors
Commercially available network encryptors from Zybersafe [16] that support the 100G Ethernet line interface and implement AES in the 256-Galois counter mode were used. The encryptors support a LAN and WAN interface for transport of the unencrypted and encrypted data, respectively. The keys were refreshed every 2 minutes or after a maximum of $2^{32}$ message frames, with the frame size being 864 bits.
The secret key material generated at the end of the QKD protocol was offloaded to a file, and a simple key management server running on each workstation loaded the content of these files into the encryptor using a serial connection. A pair of AES 256-bit keys (512 bits in total) at a time were supplied to the respective Ethernet encryptor for both encryption and decryption of the messages. Key IDs for these pairs were sequentially generated and exchanged on the encryption channel to verify that both encryptors used the correct key.
III Implementation & Results
The goal of any QKD protocol implementation is to maximize the SKF under the given channel conditions. Before delving into the general details of the protocol runs, we first describe our efforts to overcome the challenge of polarization optimization, which plays a crucial role in the long-term and autonomous operation of qTReX as well as in maximisation of the SKF.
III.1 Dynamic polarization optimization

### v2

# Future proofing network encryption technology with continuous-variable quantum key distribution

@@
Future proofing network encryption technology with continuous-variable quantum key distribution
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
@@
We demonstrate a proof-of-concept establishment of quantum-secure data transfer links in field trials at two locations in Denmark: on the campus of Technical University of Denmark in Lyngby and between power grid nodes owned and operated by Energinet in Odense. Several different links, implemented physically using optical ground wires, underground fibers as well as their combinations, were investigated. Coherent ‘quantum’ states at 1550 nm, prepared and measured using a semi-autonomous continuous-variable quantum key distribution (CVQKD) prototype, were multiplexed in wavelength with ‘classical’ 100Gbps encrypted data traffic from a pair of commercial layer-2 network encryption devices operating at around 1300 nm. Under the assumptions of real-time data processing, we estimate average secret key rates of $434.8,148.6$, and $78.3$ kbps in the asymptotic limit for diverse channels with losses (at 1550 nm) of 4.1, 5.5, and 6.7 dB, respectively. The demonstrations permit an evaluation of the prototype’s tolerance to harsh field conditions and showcase that CVQKD can serve as an additional layer to protect sensitive network traffic propagating on insecure channels.
1 Introduction
@@
The explosion in the data traffic volumes in the last few years, driven by rapid growth of data centers and cloud service providers, poses significant security challenges [1]. This is due to the nature of the data involved, which ranges from critical national infrastructure and user identities to financial secrets such as intellectual property, all of which are at risk. In particular, the energy infrastructure has been increasingly targeted, as evidenced by incidents such as the breach at a nuclear power plant in India [2], the Colonial Pipeline ransomware case in USA [3], and most recently, the coordinated hacking of 22 different energy companies in Denmark [4]. These events highlight a worrisome trend of increasing frequency, sophistication, and involvement of nation-states in cyberattacks.
@@
While implementing good practices such as timely application of patches and measures to prevent exploitation of vulnerabilities can never be underscored more, cryptographic solutions such as encryption add a protective layer on the sensitive data to foil adversaries. In the industrial context, a typical approach involves the use of hardware encryption devices, also known as network encryptors or Ethernet encryptors [5]. The devices, employing symmetric cryptographic algorithms such as advanced encryption standard (AES), operate in pairs, utilizing the same sequence of pre-configured keys for encryption and decryption of data.
@@
One of the most secure methods to distribute random symmetric keys across an insecure communication channel is provided by quantum key distribution (QKD) [6, 7, 8]. In QKD, two users, Alice and Bob, can share quantum correlations through the exchange of quantum-optical signals, while ensuring that an eavesdropper, Eve, cannot intercept these correlations without getting disclosed. This is because any eavesdropping attempt introduces noise and/or loss that Alice and Bob can quantitatively estimate, assuming pre-calibrated devices.
@@
In this work, we showcase an example of quantum-classical integration for securing data in transit using QKD and network encryption. A semi-autonomous, telecom-rack mountable continuous-variable (CV) QKD system [14, 15, 16, 17], operating at 1550 nm, produced a secret keystream buffer for a pair of layer 2 network encryptors [18], communicating simultaneously over the same channel using 100G optical transceivers at 1300 $\pm$ 10 nm. Demonstrations were conducted at two locations in Denmark with different combinations of underground and aerial fibers acting as the channel for the propagation of the C-band quantum states and the O-band optical signals, with off-the-shelf coarse wavelength division multiplexers (CWDMs) used for their union and separation.
Table 1 summarizes the lengths and losses (at 1550 nm) of the deployed channels, labelled L1-L4.
@@
Table 1: Information about the links investigated in this work and the secret key rate (SKR) obtained thereon. All links were loopbacks in the sense that both the input and output fiber were physically accessible at the same place. The loss and length at Lyngby were estimated using simple optical power and pulse delay measurements. At Odense, these quantities were determined using an optical time domain reflectometer.
Link label Location Channel (fiber) type Length, km Physical loss, dB SKR, kbps
L1 Lyngby Underground only $0.55$ $5.5$ $148.6$
L2 Odense Aerial only $13.2$ $4.1$ $434.8$
L3 Odense Underground + Aerial $15.3$ $6.7$ $78.3$
L4 Odense Aerial only $25.9$ $8.9$ $0.0$
The entire quantum-secure communication scheme, illustrated in Figure 1, was first implemented on L1, a loopback single-mode fiber (SMF) under the Lyngby campus of Technical University of Denmark (DTU).
Figure 1: Practical network encryption with continuous-variable quantum key distribution. (Top row) Map of Denmark in the center, with the geographical locations for the demonstrations indicated using the cyan and yellow pins. On the left is a satellite view showing the power substation locations in Odense, and on the right is a zoomed-in map of DTU. (Center row) Optical connections are depicted using dashed-green, dotted-blue or solid-orange lines depending on the wavelength(s) of light traversing through them, with a single/pair of line(s) indicating simplex/duplex operation. A fixed attenuator and a pair of patchcords provide a path for signals at $\sim$ 1300 nm to traverse from RX to TX. (Bottom row) The CVQKD-TX and RX modules comprise a solid plate with the optical/optoelectronic components on one side (as shown) and the necessary control electronics on the other. The main text contains a detailed description of all relevant devices as well as the system operation. TX: transmitter, RX: receiver, DPI: digital processing interface.
This was followed by temporary installation and operation of the QKD system at the premises of Energinet in Odense. Here, we explored three different loopback topologies L2-L4 using optical ground wires (OPGWs) running between power substations located at Fraugde and South East Odense (Odense SØ), as indicated in the top-left map in Fig. 1. More specifically, two of the links (L2 and L4) were a 2$\times$ and 4$\times$ back-and-forth loop of these OPGWs / aerial fibers, while L3 was a combination of an OPGW and an underground SMF.
The last column in Table 1 shows the final results—secret key rate (SKR) estimates in the asymptotic limit—obtained on these links, assuming real-time signal and data processing at the operating rate ($B=20\,$MBaud) of our CVQKD system. In practice, keys were generated offline. The large wavelength separation (1550 vs 1300 nm) of the quantum and classical signals together with the usage of balanced coherent detection ensured that despite being more than 7 orders of magnitude higher in launch power, the classical signal did not have any observable influence on the quantum signal noise characteristics. The system’s performance was challenged to some extent by the harsh environmental conditions (especially for L2-L4) which made it harder to control the channel parameters. For instance, on L4, the high untrusted loss as well as relatively high untrusted/excess noise prevented achieving a positive SKR. Nevertheless, on L1-L3, where keys could be generated successfully, the SKR values are sufficiently large to support the refresh rate requirement of 30 AES 256-bit secret key pairs per hour for continuous and real-time operation of the network encryptors. Our work thus serves as a proof-of-concept demonstration of using QKD to ‘future proof’ network encryption technology. We note that QKD has been previously used in a deployed electric utility fiber network for secure authentication of smart grid communications [19].
@@
The paper is organized as follows: In the next section, we provide details of the complete quantum-classical setup that enables future proofing network encryption with CVQKD. This is followed by a description of the conducted experiments and the obtained results. Thereafter, we discuss the results, in particular, focusing on the channel fluctuations and their influence on the live optimization of the state of polarization (SOP) of the 1550 nm signal. Finally, we provide an outlook before concluding the paper.
2 Scheme: Signals, Channels & Devices
At the heart of the scheme shown in Fig. 1 is the CVQKD system comprising a transmitter (TX) and receiver (RX), implemented as standard 3U 19-inch modules suitable for telecom rack installation. Coined qTReX [16], this system uses randomness from a state-of-the-art quantum random number generator (QRNG) [20] and a host of digital signal processing (DSP) methods [21, 14, 17] to enable the sharing of quantum correlations. Classical information processing steps, applied offline on sequences of data frames from the CVQKD-TX and -RX, utilize these correlations to distill secret keys for the encryptors.
Currently, the entire setup can be operated with minimal user intervention. Once the classical data processing stack is also brought online, the entire operation from quantum states’ production to live data encryption would become fully autonomous. With a few hardware modifications, the qTReX system can be adopted in data-center-like environments [22].
@@ 2 Scheme: Signals, Channels & Devices
In this section, we present the building blocks of the setup. More specifically, we elaborate on the different communication channels, the exchanged signals, the connected devices, and their roles in the different stages of the QKD protocol. For an introduction to the principles of CVQKD as well as typical CVQKD schemes, we refer the reader to reviews in Refs. [7, 23, 24, 25].
2.1 Signals
@@
However, the quantum-optical signal coded at the sideband frequencies is relatively dim. To address the tasks of carrier recovery and clock recovery—necessary for phase-reference alignment and synchronization of the RX and TX—the CVQKD system utilizes some ancillary signals with higher signal-to-noise ratio (SNR). These ancillary signals and the quantum-optical signal (multiplexed in time and frequency; as explained later) were the only signal connecting the CVQKD-TX and RX. To emphasize, neither any kind of electronic triggering signal nor any reference-clock-sharing between the RX and TX was implemented.
2.2 Quantum channel
[…]
2.3 Classical channel
While the CVQKD devices were connected solely through the quantum channel, their respective digital processing interfaces (DPIs; described in subsection 2.7) were connected ‘classically’. Specifically, a pair of workstations on the local network formed a classical channel (not shown in Fig. 1) for the exchange of data between TX and RX during the classical phase of the QKD protocol. Ideally, this classical channel should be authenticated [7], typically achieved using a part of the secret key from an earlier round of the protocol. However, to simplify the overall scheme, we did not implement such authentication.
2.4 Encryption channel
For high-speed (1-100 Gbps) network encryption over dedicated point-to-point links serving data centers, cloud service providers, and large enterprises, fiber-optic small form factor pluggables (SFPs) are the standard solution for distances spanning several kilometers [27].
We used off-the-shelf 100G QSFP28 4WDM-40 hot-pluggable modules [28] for transmitting and receiving the encrypted data. As depicted by the dashed-green line pairs in Fig. 1, these transceiver modules operate in duplex mode. To ensure bidirectional traffic at 1300 $\pm$ 10 nm without saturating the SFP receiver at the encryptor on the TX side, we inserted a fixed attenuator (loss $\sim$ 5 dB at 1300 nm) between the CWDMs. The encryption channel thus constituted the quantum channel and fiber patch cords with the fixed attenuator (TX to RX: solid-orange line, RX to TX: dashed-green line).
2.5 CVQKD-TX
@@
The flowchart in Fig. 2(a) illustrates the DSP steps for creating the digital input to the DAC, while the spectrum of the complex waveform, whose real (imaginary) part is related to RF1 (RF2), is depicted in Fig. 2(b).
Figure 2: Digital signal processing (DSP) in qTReX. (a) Flowchart showing the DSP procedure for quantum state encoding in a $B=20\,$MHz wide sideband and multiplexing of ancillary signals in time and frequency. (b) Spectrum of the complex signal, the real and imaginary parts of which are played out using a two-channel DAC. Due to time multiplexing, the quantum signal and CAZAC sequence may share the same frequency band, however, the former is $\sim 50$ times lower in power than the latter. (c) Spectra of the ADC captured signals under various configurations of the optical switches in CVQKD-RX, detailed further in the text. Vac: vacuum state, Q/PRNG: quantum/pseudo random number generator. CAZAC: constant amplitude zero auto correlation.
Uniformly distributed random numbers from the QRNG were transformed into discrete Gaussian-distributed numbers with 6 bits of resolution and a range of 7 standard deviations [20]. These formed the amplitudes and phases of the quantum states, and to these ‘quantum’ symbols, a constant amplitude zero auto correlation sequence was multiplexed in time, while a quadrature phase shift keying (QPSK) symbol sequence containing a known header and frame identification (ID) information was multiplexed in frequency [17]. The subsequent step of pulse shaping was performed by upsampling to the DAC sampling rate of 1 GSamples/s and applying root raise cosine filtering [14, 15]. After digital single sideband modulation or complex upconversion of the two 20 MHz wide sidebands at 80 and 160 MHz, two pilot tones at 25 MHz and 120 MHz were also multiplexed in frequency as depicted in Fig. 2(b).
Referring to Fig. 1 (bottom), the optical power at the output of the IQ modulator was monitored using a 1:99 fiber splitter and power meter. We could independently set the 1) modulation strength associated with the quantum states, 2) power of the pilot tones, and 3) power in the QPSK signal band to desired levels through a combined digital control of signal amplitudes during the DSP procedure, tuning of the voltage output levels of the DAC, and the optical attenuation using an electronically controlled variable optical attenuator after the IQ modulator.
All fiber-optic components featured polarization maintaining fiber (PMF) pigtails, except the SMF-connectorized wavelength filter and isolator (added to prevent Trojan-horse attacks [30]). The signal $\text{TX}_{\text{out}}$ was fed by a simplex cable to the input of CWDM-TX.
2.6 CVQKD-RX
@@
Due to arbitrary polarization fluctuations in the channel, the incoming signal $\text{RX}_{\text{in}}$ required dynamic polarization control; section 3.1 describes the details of this procedure. The signal polarization is in a desired state when it aligns with the polarization of the LLO, where after, the ADC captured the modulated signal.
@@
2.7 Digital Processing Interfaces
Two workstations equipped with high-end graphic processing unit (GPU) cards formed the DPIs for operating the entire QKD protocol. PCI Express based DAC and ADC cards, housed in these workstations, were used for driving the modulator in CVQKD-TX and acquiring the detector output in CVQKD-RX, respectively, as explained above. The DSP cycles running on each of them allowed the quantum symbols encoded at CVQKD-TX to be eventually decoded at CVQKD-RX, i.e., fulfil the state preparation and measurement tasks. Figure 3 in Ref. [17] shows a flowchart of the DSP running on CVQKD-RX, complementary to the one Fig. 2(a).
Thereafter, the workstations communicated over the classical channel to perform the remaining data processing stages of the protocol, namely information reconciliation (IR), parameter estimation (PE), and privacy amplification (PA). Usage of GPUs speeds up the execution of both IR and PA [31, 32].
2.8 Network encryptors
Commercially available network encryptors from Zybersafe [18] that support the 100G Ethernet line interface and implement AES in the 256-Galois counter mode were used. The encryptors support a LAN and WAN interface for transport of the unencrypted and encrypted data, respectively. The keys are refreshed every 2 minutes or after a maximum of $2^{32}$ messages, with message size being 864 bits.
The secret key material generated at the end of the QKD protocol was offloaded to a file, and a simple key management server running on each workstation loaded the content of these files into the encryptor using a serial connection. A pair of AES 256-bit keys (512 bits in total) at a time were supplied to the respective Ethernet encryptor for both encryption and decryption of the messages. Key IDs for these pairs were sequentially generated and exchanged on the encryption channel to ascertain that both encryptors used the correct key.
3 Implementation & Results
The goal of any QKD protocol implementation is to maximize the SKR under the given channel conditions. Before delving into the general details of the protocol runs, we first describe our efforts to overcome the challenge of polarization optimization, which plays a crucial role in the long-term and autonomous operation of qTReX as well as in maximisation of the SKR.
3.1 Dynamic polarization optimization